OpenStudy (anonymous):

in a math class of 26 students each girl drew a tringle and each boy drew a rectangle if there were 92 sides drawn by students how mant girls were in th class

OpenStudy (anonymous):

There are 3 sides to a triangle and 4 sides to a rectangle. Let g = number of girls and b = number of boys then since we have two variables, we'll need two equations. . g + b = 26 (equation 1) 3g + 4b = 92 (equation 2) . solving equation 1 for g, we get: g = 26-b . Substituting the above into equation 2 we can then solve for b: 3g + 4b = 92 3(26-b) + 4b = 92 78 - 3b + 4b = 92 78 + b = 92 b = 14 (number of boys) . Substitute the above into equation 1 and solve for g: g + b = 26 g + 14 = 26 g = 12 (number of girls)